OpenStudy (anonymous):

Solve by Factoring: 2a^2-a-21=0

OpenStudy (anonymous):

@wolf1728

OpenStudy (anonymous):

Doing the AC method, you take A times C, so 2 times -21 which equals -42. so what two numbers multiply to get -42 and add to get -1. 6 and -7 multiply to get -42 and add to get -1 so you use those two number to set up your equation. y=(x+6)(x-7) then set each part equation to 0. 0=x+6=-6 0=x-7=7 So the answer: x=-6 and x=7

OpenStudy (wolf1728):

2a^2-a-21=0 Divide by 2 a² -.5a -10.5 factors of -10.5 = 7 * 1.5 3 * 3.5 we need a negative third term and a negative second term so we change signs accordingly until we get: (a + 3) • (a-3.5)

OpenStudy (anonymous):

thanks... Do you know what What times what= 6 and add up to -2.5?

OpenStudy (wolf1728):

Hmm where are those numbers in the factoring?

OpenStudy (wolf1728):

Anyway -4 * -1.5 = 6 But do not add to -2.5 -4 * 1.5 = -6 and -4 + 1.5 = -2.5 Are you sure about the signs in that question?